A corner shelf is made in the shape of a triangle. The shelf has a total area of 324 sq. cm., with a height of 18 cm. What is the length of the base of the triangle?

Answer by Boreal(15235)

A=(1/2)bh
324=(1/2)*b*18=9b
divide by 9 both sides
b=36 cm.